# Efficient Diversity-based Experience Replay for Deep Reinforcement Learning

### v2

@@ Abstract
Experience replay is widely used to improve learning efficiency in reinforcement learning by leveraging past experiences. However, existing experience replay methods, whether based on uniform or prioritized sampling, often suffer from low efficiency, particularly in real-world scenarios with high-dimensional state spaces. To address this limitation, we propose a novel approach, Efficient Diversity-based Experience Replay (EDER). EDER employs a deterministic point process to model the diversity between samples and prioritizes replay based on the diversity between samples. To further enhance learning efficiency, we incorporate Cholesky decomposition for handling large state spaces in realistic environments. Additionally, rejection sampling is applied to select samples with higher diversity, thereby improving overall learning efficacy. Extensive experiments are conducted on robotic manipulation tasks in MuJoCo, Atari games, and realistic indoor environments in Habitat. The results demonstrate that our approach not only significantly improves learning efficiency but also achieves superior performance in high-dimensional, realistic environments.
@@ 1 Introduction
In recent years, Deep Reinforcement Learning Arulkumaran et al. (2017); Wang et al. (2024, 2025) has surged in popularity, achieving remarkable success in complex decision-making tasks. DRL has been successfully applied to games Schrittwieser et al. (2020); Silver et al. (2017); Chen et al. (2024), robotic control Andrychowicz et al. (2020); Levine et al. (2016); Todorov et al. (2012), autonomous driving Kiran et al. (2021); Yang et al. (2023b, 2024) , and other domains, demonstrating its powerful learning and decision-making capabilities.
@@
Figure 1: Sample distribution comparison the replay buffer. Left: Uniform sampling results in an imbalanced distribution, with some data types overrepresented and others underrepresented. Right: Our method achieves a more balanced and diverse selection of samples, enhancing overall diversity and improving learning efficiency.
@@
However, DRL still faces significant challenges in practical applications, particularly in handling sparse reward signals Hare (2019); Vecerik et al. (2017), high-dimensional state spaces Warnell et al. (2018); Ibrahimi et al. (2012), and low sample efficiency. Sparse reward signals make it difficult for agents to learn effective policies from limited positive feedback, resulting in slow and inefficient learning processes. Additionally, high-dimensional state spaces further complicate the learning process and increase computational burdens, making existing methods inefficient in large-scale and complex environments.
@@
To address these issues, Experience Replay (ER) has been widely adopted as a key mechanism. ER improves sample efficiency and stabilizes the learning process by storing the agent’s past experiences and randomly sampling from them for training. Despite the improvements ER offers in sample efficiency, existing methods still suffer from inefficiency and suboptimal performance in high-dimensional state spaces. Recent studies Andrychowicz et al. (2020); Levine et al. (2016); Todorov et al. (2012); Jiang et al. (2024); Yang et al. (2023a); Zhao and Tresp (2018); Fang et al. (2019) have focused on enhancing ER’s sampling strategies to improve their applicability and efficiency in complex environments. For instance, Hindsight Experience Replay (HER) Andrychowicz et al. (2017) generates more positive feedback samples to enhance learning efficiency; Prioritized Experience Replay (PER) Schaul et al. (2015) assigns priorities to samples based on their temporal difference (TD) errors; and Topological Experience Replay (TER) Hong et al. (2022) builds a trajectory graph and performs breadth-first updates from terminal states. Large Batch Experience Replay (LaBER) Lahire et al. (2022) improves sample efficiency by sampling large batches and performing focused updates. The Reducible Loss (ReLo) method Sujit et al. (2023) ranks samples based on their learnability, measured by consistent loss reduction. However, these approaches generally struggle to efficiently select valuable samples in high-dimensional state spaces, leading to persistent issues of low efficiency and high-dimensional state space challenges in DRL.
@@
To tackle these challenges, we propose a novel Experience Replay framework, Efficient Diversity-based Experience Replay (EDER). EDER utilizes Determinantal Point Processes (DPP) Kulesza et al. (2012) to model the diversity among samples and determines replay priorities based on this diversity, effectively avoiding the redundant sampling of ineffective data points. Furthermore, to handle high-dimensional state spaces in real-world environments, EDER employs Cholesky decomposition Krishnamoorthy and Menon (2013), significantly reducing computational complexity. Combined with rejection sampling techniques Neal (2003), EDER selects samples with higher diversity for training, thereby further enhancing overall learning efficiency.
@@
Our main contributions are as follows. Firstly, we propose the Efficient Diversity-based Experience Replay (EDER) framework, which prioritizes sample diversity and significantly enhances experience replay (ER) efficiency, especially in high-dimensional state spaces and environments with sparse rewards. Secondly, we introduce Cholesky decomposition and rejection sampling to effectively address computational bottlenecks in large state spaces and optimize the ER mechanism by selecting more diverse samples. Lastly, we conduct extensive experimental validations across multiple complex environments, including Habitat Savva et al. (2019); Szot et al. (2021), Atari games Mnih (2013), and MuJoCo Todorov et al. (2012). The results demonstrate that EDER not only significantly improves learning efficiency but also achieves superior performance in high-dimensional, realistic environments, thereby validating its effectiveness and adaptability in various complex settings.
@@ 2 Preliminaries
where $L_{Y}$ is the principal submatrix of $L$ indexed by the elements in $Y$, and $I$ is the identity matrix. The determinant $\text{det}(L_{Y})$ measures the diversity of $Y$ by the volume spanned by the vectors associated with $Y$. In practice, the kernel matrix $L$ is often the Gram matrix $L=X^{T}X$, where each column of $X$ represents a feature vector of an element in $Y$. The geometric interpretation of DPPs implies that subsets with more orthogonal feature vectors—indicating higher diversity—are more likely to be selected. This makes DPPs effective for sampling diverse trajectories and goals in reinforcement learning, where diversity in the experience buffer is crucial for robust learning.
@@ 3.1 Diversity-Based Trajectory Selection Module
Trajectory Segmentation. The entire sequence of state transitions during an interaction, denoted as $\tau$, is segmented into several partial trajectories $\tau_{j}$ of length $b$. Each segment $\tau_{j}$ covers transitions from state $s_{n}$ to $s_{n+b-1}$, allowing for detailed capture of dynamics between state transitions. In this part, with a sliding window of $b=2$, a trajectory $\tau$ can be divided into $N_{p}$ segments of partial trajectories.
$\tau_{i}=\bigl{\{}\{\underbrace{s_{0},s_{1}}_{\tau_{1}}\},\{\underbrace{s_{2},s_{3}}_{\tau_{2}}\},\{\underbrace{s_{4},s_{5}}_{\tau_{2}}\},\ldots,\{\underbrace{s_{T-1},s_{T}}_{\tau_{N_{p}}}\}\bigr{\}}$
@@
Here, $L_{\tau_{j}}$ is the kernel matrix constructed from the state transitions within trajectory $\tau_{j}$, defined as:
@@
Equation (4) provides a comprehensive measure, accurately reflecting the overall diversity of the trajectory. We employ a non-uniform sampling strategy to prioritize trajectories with higher diversity:
@@
where $N_{e}$ is the total number of trajectories in the replay buffer. This strategy enhances learning efficiency by increasing the likelihood of selecting highly diverse trajectories, thereby enabling the agent to effectively learn and adapt to various environmental conditions.
Although the determinant effectively measures diversity, its direct computation in high-dimensional state spaces is computationally intensive, especially for large trajectory lengths $b$. Therefore, we employs Cholesky decomposition and rejection sampling to optimize computation speed in the following section. The diversity metric $d_{\tau_{j}}$ quantifies the independence and diversity of state vectors within trajectory $\tau_{j}$ using the determinant.
@@ 3.2 Improving Computational Efficiency
After Cholesky decomposition, the time complexity of determinant calculation for each segment has been reduced from $O(n^{3})$ to $O(n^{2})$.
Rejection Sampling. Numerous trajectories are utilized in training, resulting in high sampling inefficiency. To enhance sampling efficiency, we introduce rejection sampling. This method effectively samples trajectory segments in high-dimensional state spaces, where direct sampling incurs significant computational overhead. By prioritizing segments with higher diversity scores, rejection sampling minimizes the selection of less informative segments. Consequently, computational resources are focused on the most diverse and relevant experiences, reducing overhead and ensuring that the replay buffer contains the most valuable transitions. The rejection sampling process is detailed as follows: First, for each trajectory segment $\tau_{j}$, we compute its diversity score $Q_{j}$ using Equation (2) and (7) :
@@
Next, we determine a bounding constant $M$ to ensure that for all trajectory segments $\tau_{j}$, the inequality $Q_{j}\leq\frac{M}{N}$ holds, where $N$ is the total number of trajectory segments in the replay buffer. We achieve this by selecting:
$M=N\cdot\max Q_{j}$ (9)
The choice of $M$ guarantees that the acceptance probability remains within the valid range $[0,1]$. During the sampling process, we uniformly sample a candidate trajectory segment $\tau^{\prime}$ from the proposal distribution $q(\tau^{\prime})=\frac{1}{N}$. Subsequently, we generate a uniformly distributed random number $u\sim U(0,1)$. If the sampled number satisfies:
$u\leq\alpha=\frac{d_{\tau^{\prime}}}{\max Q_{j}},$ (10)
we accept the candidate trajectory segment $\tau^{\prime}$; otherwise, we reject it and resample. By iteratively executing the sampling procedure, we acquire the required number $m$ of diverse trajectories to update the policy.
@@ 3.3 Time Complexity Analysis
Proof is in Appendix A. Based on Theorem 2, the integration of Cholesky decomposition and rejection sampling significantly reduces the overall computational complexity of the EDER algorithm, especially when dealing with large segment lengths b. This improvement makes the EDER algorithm more efficient and scalable for high-dimensional reinforcement learning tasks, enhancing its applicability in complex environments.
@@
15:     for each $\tau_{j}$ in $Q$ do
16:        Calculate acceptance $\alpha=\frac{d_{\tau^{\prime}}}{\max Q_{j}}$
17:        Accept $\tau_{j}$ if $u\leq\alpha$, else resample
@@
19:     Sample experiences to update $\mathcal{D}$
20:     Optimize $\theta$ with sampled experiences
@@ 4 Experiments
Our experiments aim to rigorously evaluate the performance of the proposed Efficient Diversity-based Experience Replay (EDER) method across multiple environments, focusing on its effectiveness compared to established baseline methods. The experiments are conducted in Mujoco, Atari, and real-life Habitat environments, each selected to highlight different aspects of EDER’s capabilities. Detailed environment settings are provided in Appendix  C.
@@ 4.4 Ablation Studies
In our ablation studies, we evaluate the impact of rejection sampling and Cholesky decomposition on the training efficiency of EDER. As shown in Figure 5, omitting rejection sampling reduces exploration efficiency, slows convergence, and lowers success rates. In contrast, excluding Cholesky decomposition leads to increased instability and longer training times, ultimately decreasing learning efficiency. Furthermore, when comparing the training times of the HER-based algorithm, EDER demonstrates competitive performance, showcasing its computational efficiency. Additional experiments varying $m$ (the number of diverse trajectories) and $b$ (the trajectory length) are provided in Appendix B. Our results show that increasing $m$ improves exploration by selecting more diverse trajectories, thus helping to avoid local optima. However, overly large $m$ increases computational cost without yielding proportional performance gains, which results in slower convergence. Similarly, longer trajectories $b$ offer richer contextual information for learning and improve exploration in temporally dependent environments. However, excessively long trajectories can lead to instability, impeding training efficiency.
Figure 5: Training times for EDER and baselines on Push task. EDER shows competitive times with Rejection Sampling (R.S.) and Cholesky Decomposition (C.D.).
@@ 5 Related Work
Prioritized Experience Replay (PER) Schaul et al. (2015) enhances learning by focusing on high TD-error samples, prioritizing informative experiences. Various extensions to PER have been proposed, such as the actor-critic-based PER Saglam et al. (2022), which dynamically adjusts sampling priorities to balance exploration and exploitation; Attentive PER Sun et al. (2020) uses attention mechanisms to replay experiences relevant to the current learning phase, enhancing training efficiency. Additionally, recent studies have introduced new priority criteria to enhance PER’s effectiveness. Relo Sujit et al. (2023) define the learnability of transitions as a priority criterion, prioritizing samples that consistently reduce training loss.TER Hong et al. (2022) builds a trajectory graph and prioritizes updates breadth-first from terminal states; LaBER Lahire et al. (2022) enhances efficiency by leveraging large batch sampling with focused updates. FSER Yu et al. (2024) combines frequency and similarity indices to prioritize rare and policy-aligned experiences.  Wei et al. (2021) integrates transition revisit frequency with TD error for more effective replay buffer prioritization.

### v4

@@ Abstract
Experience replay is widely used to improve learning efficiency in reinforcement learning by leveraging past experiences. However, existing experience replay methods, whether based on uniform or prioritized sampling, often suffer from low efficiency, particularly in real-world scenarios with high-dimensional state spaces. To address this limitation, we propose a novel approach, Efficient Diversity-based Experience Replay (EDER). EDER employs a determinantal point process to model the diversity between samples and prioritizes replay based on the diversity between samples. To further enhance learning efficiency, we incorporate Cholesky decomposition for handling large state spaces in realistic environments. Additionally, rejection sampling is applied to select samples with higher diversity, thereby improving overall learning efficacy. Extensive experiments are conducted on robotic manipulation tasks in MuJoCo, Atari games, and realistic indoor environments in Habitat. The results demonstrate that our approach not only significantly improves learning efficiency but also achieves superior performance in high-dimensional, realistic environments.
@@ 1 Introduction
In recent years, Deep Reinforcement Learning François-Lavet et al. (2018); Wang et al. (2024a) has surged in popularity, achieving remarkable success in complex decision-making tasks. DRL has been successfully applied to games Schrittwieser et al. (2020); Silver et al. (2017), robotic control Andrychowicz et al. (2020); Levine et al. (2016), autonomous driving scenarios including traffic light control Yang et al. (2023b, a, 2024), and other domains, demonstrating its powerful learning and decision-making capabilities.
@@
Figure 1: Sample distribution comparison of the replay buffer. Left: Uniform sampling results in an imbalanced distribution, with some data types overrepresented and others underrepresented. Right: Our method achieves a more balanced and diverse selection of samples, enhancing overall diversity and improving learning efficiency.
@@
However, DRL still faces significant challenges in practical applications, particularly in handling sparse reward signals Hare (2019), high-dimensional state spaces Ibrahimi et al. (2012), and low sample efficiency Yarats et al. (2021); Wang et al. (2024b). Sparse reward signals make it difficult for agents to learn effective policies from limited positive feedback, resulting in slow and inefficient learning processes. Additionally, high-dimensional state spaces further complicate the learning process and increase computational burdens, making existing methods inefficient in large-scale and complex environments.
@@
To address these issues, Experience Replay (ER) has been widely adopted as a key mechanism. ER improves sample efficiency and stabilizes the learning process by storing the agents’ past experiences and randomly sampling them for training. Despite the improvements ER offers in sample efficiency, existing methods still suffer from inefficiency and suboptimal performance in high-dimensional state spaces. Recent studies Andrychowicz et al. (2020); Levine et al. (2016); Todorov et al. (2012); Jiang et al. (2024); Zhao and Tresp (2018); Fang et al. (2019) have focused on enhancing ER’s sampling strategies to improve their applicability and efficiency in complex environments. For instance, Hindsight Experience Replay (HER) Andrychowicz et al. (2017) generates more positive feedback samples to enhance learning efficiency; Prioritized Experience Replay (PER) Schaul et al. (2015) assigns priorities to samples based on their temporal difference (TD) errors; and Topological Experience Replay (TER) Hong et al. (2022) builds a trajectory graph and performs breadth-first updates from terminal states. Large Batch Experience Replay (LaBER) Lahire et al. (2022) improves sample efficiency by sampling large batches and performing focused updates. The Reducible Loss (ReLo) method Sujit et al. (2023) ranks samples based on their learnability, measured by consistent loss reduction. However, these approaches generally struggle to efficiently select valuable samples in high-dimensional state spaces, leading to persistent issues of low efficiency and high-dimensional state space challenges in DRL.
@@
To tackle these challenges, we propose a novel Experience Replay framework, Efficient Diversity-based Experience Replay (EDER). EDER utilizes Determinantal Point Processes (DPP) Kulesza et al. (2012) to model the diversity among samples and determines replay priorities based on this diversity, effectively avoiding the redundant sampling of ineffective data points. Furthermore, to handle high-dimensional state spaces in real-world environments, EDER employs Cholesky decomposition Krishnamoorthy and Menon (2013), significantly reducing computational complexity. Combined with rejection sampling techniques Neal (2003); Azadi et al. (2018), EDER selects samples with higher diversity for training, thereby further enhancing overall learning efficiency.
@@
Our main contributions are as follows. Firstly, we propose the Efficient Diversity-based Experience Replay (EDER) framework, which prioritizes sample diversity and significantly enhances experience replay (ER) efficiency, especially in high-dimensional state spaces and environments with sparse rewards. Secondly, we introduce Cholesky decomposition and rejection sampling to effectively address computational bottlenecks in large state spaces and optimize the ER mechanism by selecting more diverse samples. Lastly, we conduct extensive experimental validations across multiple complex environments, including Habitat Savva et al. (2019); Puig et al. (2023), Atari games Mnih (2013), and MuJoCo Todorov et al. (2012). The results demonstrate that EDER not only significantly improves learning efficiency but also achieves superior performance in high-dimensional, realistic environments, thereby validating its effectiveness and adaptability in various complex settings.
@@ 2 Preliminaries
where $L_{Y}$ is the principal submatrix of $L$ indexed by the elements in $Y$, and $I$ is the identity matrix. The determinant $\text{det}(L_{Y})$ measures the diversity of $Y$ by the volume spanned by the vectors associated with $Y$. In practice, the kernel matrix $L$ is often the Gram matrix $L=X^{T}X$, where each column of $X$ represents a feature vector of an element in $Y$. The geometric interpretation of DPPs implies that subsets with more orthogonal feature vectors—indicating higher diversity—are more likely to be selected. This makes DPP effective for sampling diverse trajectories and goals in reinforcement learning, where diversity in the experience buffer is crucial for robust learning.
@@ 3.1 Diversity-Based Trajectory Selection Module
Trajectory Segmentation. The entire sequence of state transitions during an interaction, denoted as $\tau$, is segmented into several partial trajectories $\tau_{j}$ of length $b$. Each segment $\tau_{j}$ covers transitions from state $s_{n}$ to $s_{n+b-1}$, allowing for detailed capture of dynamics between state transitions. For clarity, we set a sliding window of $b=2$ in this part, while other values are explored in the ablation studies. Under this setting, a trajectory $\tau$ can be divided into $N_{p}$ partial segments.
$\tau=\bigl{\{}\{\underbrace{s_{0},s_{1}}_{\tau_{1}}\},\{\underbrace{s_{2},s_{3}}_{\tau_{2}}\},\{\underbrace{s_{4},s_{5}}_{\tau_{3}}\},\ldots,\{\underbrace{s_{T-1},s_{T}}_{\tau_{N_{p}}}\}\bigr{\}}$
@@
Intuitively, the determinant quantifies the $n$-dimensional volume spanned by the embedded state transitions in $\tau_{j}$, assigning higher values to sets of transitions that are more linearly orthogonal and thus more diverse. Here, $L_{\tau_{j}}$ is the kernel matrix constructed from the state transitions within trajectory $\tau_{j}$, defined as:
@@
Equation (4) provides a comprehensive measure, effectively reflecting the overall diversity of the trajectory. We employ a non-uniform sampling strategy to prioritize trajectories with higher diversity:
@@
where $N_{e}$ is the total number of trajectories in the replay buffer, this strategy enhances learning efficiency by increasing the likelihood of selecting highly diverse trajectories, thereby enabling the agent to effectively learn and adapt to various environmental conditions.
Although the determinant effectively measures diversity, its direct computation in high-dimensional state spaces is computationally intensive, especially for large trajectory lengths $b$. Therefore, we employ Cholesky decomposition and rejection sampling to optimize computation speed in the following section. The diversity metric $d_{\tau_{j}}$ quantifies the independence and diversity of state vectors within trajectory $\tau_{j}$ using the determinant.
@@ 3.2 Improving Computational Efficiency
Here, $l_{ii}$ denotes the $i$-th diagonal element. With Cholesky decomposition, the time complexity of determinant computation for each segment is reduced from $O(n^{3})$ to $O(n^{2})$.
Rejection Sampling. Numerous trajectories are utilized in training, resulting in high sampling inefficiency. To enhance sampling efficiency, we introduce rejection sampling. This method effectively filters trajectory segments before insertion into the replay buffer, which is particularly useful in high-dimensional state spaces where storing redundant segments incurs significant computational overhead. By prioritizing segments with higher diversity scores, rejection sampling minimizes the retention of less informative segments. Consequently, computational resources are focused on the most diverse and relevant experiences, ensuring that the replay buffer contains the most valuable transitions.
The rejection sampling process is detailed as follows: First, for each trajectory segment $\tau_{j}$, we compute its diversity score $Q_{j}$ using Equation (2) and (7):
@@
Next, we determine a normalization constant $M$, defined as:
$M=\max(Q)$ (9)
This ensures that for all trajectory segments $\tau_{j}$, the acceptance probability $\alpha=\frac{Q_{j}}{M}$ remains in the valid range $[0,1]$. During the rejection sampling process, we uniformly select a candidate segment $\tau^{\prime}$ from the current batch of generated segments. Then we draw a uniformly distributed random number $u\sim U(0,1)$. If the sampled number satisfies:
$u\leq\alpha=\frac{d_{\tau^{\prime}}}{M},$ (10)
we accept the candidate trajectory segment $\tau^{\prime}$; otherwise, we reject it and resample. This process yields a set of diverse trajectory segments, which are then inserted into the replay buffer. Subsequently, training batches are sampled from the buffer according to Equation (5).
@@ 3.3 Time Complexity Analysis
The proof is in Appendix A. Based on Theorem 2, the integration of Cholesky decomposition and rejection sampling significantly reduces the overall computational complexity of the EDER algorithm, especially when dealing with large segment lengths b. This improvement makes the EDER algorithm more efficient and scalable for high-dimensional reinforcement learning tasks, enhancing its applicability in complex environments.
@@
15:     for each $Q_{j}$ in $Q$ do
16:        Calculate acceptance $\alpha=\frac{Q_{j}}{M}$
17:        Accept corresponding $\tau_{j}$ if $u\leq\alpha$, else discard
@@
19:     Sample $\mathcal{B}\sim\mathcal{D}$ using Eq. (5)
20:     Update $\theta$ using $\mathcal{B}$
@@ 4 Experiments
Our experiments aim to rigorously evaluate the performance of the proposed Efficient Diversity-based Experience Replay (EDER) method across multiple environments, focusing on its effectiveness compared to established baseline methods. The experiments are conducted in Mujoco, Atari, and real-life Habitat environments, each selected to highlight different aspects of EDER’s capabilities. Detailed environment settings are provided in Appendix C. Details are available at https://arxiv.org/abs/2410.20487.
@@ 4.4 Ablation Studies
In our ablation studies, we evaluate the impact of Rejection Sampling (R.S.) and Cholesky Decomposition (C.D.) on the training efficiency of EDER. As shown in Table 3, removing R.S. reduces exploration efficiency, slows convergence, and lowers success rates. In contrast, excluding C.D. leads to increased instability and longer training times, ultimately diminishing learning efficiency. Moreover, when comparing training times with HER-based baselines, EDER maintains competitive performance, especially in more complex tasks like PickAndPlace, demonstrating favorable trade-offs between computational cost and performance. Additional experiments varying $m$ (the number of diverse trajectories) and $b$ (the trajectory length) are provided in Appendix B. Our results indicate that increasing $m$ promotes exploration by sampling more diverse trajectories, helping to avoid local optima. However, excessive $m$ increases computational burden without proportionate performance gains, leading to slower convergence. Likewise, longer segments ($b$) provide richer context and improve exploration in temporally dependent environments, but overly long trajectories may introduce instability and reduce overall training efficiency.
PickAndPlace Task (Training Time in Minutes)
Method Time Method Time
DDPG + HER (Buffer) 80.7 DDPG + LaBER 93.6
DDPG + PER (sum-tree) 63.4 DDPG + Relo 107.1
DDPG + TER 91.9 DDPG + EDER 103.1
Push Task: DDPG + EDER Variants (in Minutes)
Method Time Method Time
EDER (b=10) 124.3 EDER (b=T) 156.0
w/o R.S. (b=10) 173.2 w/o R.S. (b=T) 182.7
w/o C.D. (b=10) 129.2 w/o C.D. (b=T) 171.3
Table 3: Training times (in minutes) for baseline methods and EDER variants on PickAndPlace and Push tasks. R.S.: Rejection Sampling; C.D.: Cholesky Decomposition.
@@ 5 Related Work
Prioritized Experience Replay (PER) Schaul et al. (2015) enhances learning by focusing on high TD-error samples and prioritizing informative experiences. Various extensions to PER have been proposed, such as the actor-critic-based PER Saglam et al. (2022), which dynamically adjusts sampling priorities to balance exploration and exploitation; Attentive PER Sun et al. (2020) uses attention mechanisms to replay experiences relevant to the current learning phase, enhancing training efficiency. Additionally, recent studies have introduced new priority criteria to enhance PER’s effectiveness. Relo Sujit et al. (2023) define the learnability of transitions as a priority criterion, prioritizing samples that consistently reduce training loss.TER Hong et al. (2022) builds a trajectory graph and prioritizes updates breadth-first from terminal states; LaBER Lahire et al. (2022) enhances efficiency by leveraging large batch sampling with focused updates. FSER Yu et al. (2024) combines frequency and similarity indices to prioritize rare and policy-aligned experiences.  Wei et al. (2021) integrates transition revisit frequency with TD error for more effective replay buffer prioritization.